# Avoid Wasted Annotation Costs in Open-set Active Learning with Pre-trained Vision-Language Model

### v1

@@ 3.3 Visual Similarity Weighting
Although CLIPN shows high performance in detecting OOD data using textual information, it encounters challenges when the number of classes increases or when the OOD data domain is similar to that of ID data. We use visual information from ID data as additional weights to address these challenges. First, the average representation $\phi_{j}$ for each class is calculated by applying the image encoder $f^{\text{img}}$ to the labeled ID data $\textbf{x}^{\text{img, ID}}$.
$\phi_{j}=\frac{1}{N_{j}^{\text{ID}}}\sum_{i=1}^{N_{j}^{\text{ID}}}{f^{\text{img}}(x_{i}^{\text{img, ID}})}.$ (5)
Next, the visual similarity weight $s_{ij}$ for each class is calculated by computing the cosine similarity between the visual representation $z_{i}^{\text{img}}$ and the average representation $\phi_{j}$, and converting it to probabilities using the softmax function.
$s_{ij}=\frac{\exp{(z_{i}^{\text{img}}\cdot\phi_{j}/\tau)}}{\sum_{k}^{K}\exp{(z_{i}^{\text{img}}\cdot\phi_{k}/\tau)}}.$ (6)
@@
$\mathbb{I}^{*}(x_{i})$ =$=\begin{cases}1,&p_{i}^{*,\text{OOD}}\geq\max_{j}{p_{ij}^{*,ID}},\\ 0,&\text{else},\end{cases}$
Figure 6: Comparison of precision changes in open-set AL based on OOD ratio. CIFAR10 and CIFAR100 are set a 40% mismatch ratio. Precision refers to the proportion of ID data in selected queries.
@@ 3.4 Self-Temperature Tuning
The temperature parameter significantly influences model confidence in predictions. The optimal temperature varies depending on the domain of the data used for OOD detection. In open-set AL, labeled ID and OOD data allow for exploring the optimal temperature parameter.
@@
$\text{Acc}(\tau)$ $=\frac{1}{N_{\text{L}}^{\text{ID}}}\sum_{i}^{N_{\text{L}}^{\text{ID}}}{\mathbb{I}(y_{i}^{\text{OOD}}=\hat{y}_{i}(\tau))},$
${\tau}^{*}$ $=\operatorname*{arg\,max}_{0<\tau}{\text{Acc}({\tau})}.$
According to Eq. 8, we apply a self-temperature tuning to explore the optimal temperature parameter by comparing the accuracy of the predicted results $\hat{\textbf{y}}^{\text{OOD}}(\tau)$ against the ground truth $\textbf{y}^{\text{OOD}}$ for ID and OOD data.
@@ 4.1 Experimental Setup
To demonstrate the efficiency of CLIPNAL, we select query strategies from standard AL and open-set AL as comparison methods. For standard AL, the following strategies are selected: (1) uncertainty-based strategy, including CONF[18], LL[21], and (2) diversity-based strategy, including CORESET[15]. For open-set AL, the chosen strategies are (1) contrastive learning-based strategy MQNet[11], (2) OOD detector-based strategy, including LfOSA[10] and EOAL[13]. To evaluate the effectiveness of query strategies, we include RANDOM, which selects data randomly without specific criteria, in the experiments. The detail of the baselines is described in Appendix. D.
Open-set Data.
@@
To construct various open-set scenarios, we consider mismatch ratios, which vary the proportion of classes in the data, and OOD ratios, which vary the proportion of OOD data in the unlabeled pool[10, 4]. The mismatch ratios are defined as 20%, 30%, and 40%. The OOD ratios are defined as 10%, 20%, 40%, and 60%. For instance, in CIFAR10 with a mismatch ratio of 20%, two classes are ID, and the remaining eight are OOD. The composition of the open-set data used in the experiments is detailed in Appendix. B. We evaluate the robustness and effectiveness of the query strategies under various scenarios. The annotation cost $C$ for AL is 500 for both datasets. The initial data for training the classifier is randomly selected from the unlabeled pool, equivalent to the defined cost. In subsequent iterations, data is selected based on the query strategy with a cost of $C$.
Figure 7: Comparison of accuracy changes in open-set AL based on OOD ratio. CIFAR10 and CIFAR100 are set a 40% mismatch ratio. The black dashed line represents the performance of a classifier trained on the entire dataset.
Figure 8: Comparison of accuracy changes in open-set AL based on mismatch ratio. CIFAR10 and CIFAR100 are set a 60% OOD ratio. The black dashed line represents the performance of a classifier trained on the entire dataset.
Implementation Details.
@@
The image classifier is ResNet18[6] in the experiments. The number of rounds $R$ for AL experiments is set to 10, with results reported for each round. Each query strategy experiment is repeated five times, with average results reported. The hyper-parameters and pre-trained models are described in Appendix. C.
Evaluation Metrics.
@@
Precision, the proportion of selected ID queries, is used to evaluate annotation cost loss in open-set AL[10]. Higher precision indicates a higher proportion of ID data in the selected queries.
@@ 4.2 Open-set Robustness
Comparison of Annotation Cost Loss.
We examine annotation cost loss through precision changes during data collection in Fig. 6. Each round’s precision indicates the proportion of ID data among the selected data at the round. RANDOM, shown in black, selects data randomly, so the proportion of OOD data in the selected queries matches the OOD ratio in the unlabeled pool. A higher precision than RANDOM indicates a higher selection of ID data than the OOD ratio.
Standard AL strategies, except CORESET, show lower precision than RANDOM in all cases, particularly as the OOD ratio increases. This indicates significant cost loss when OOD data is in the unlabeled pool. In contrast, most open-set AL strategies show higher precision than RANDOM. However, MQNet, LfOSA, and EOAL do not show consistent precision across rounds. MQNet initially focuses on purity, selecting more ID data, but later prioritizes informativeness, causing precision to drop. LfOSA and EOAL initially had low precision due to the limited OOD data for training the detection model, but they improved as more OOD data was collected. Unlike existing strategies, CLIPNAL maintains high precision throughout data collection, regardless of the OOD ratio, indicating minimal annotation cost loss.
Figure 9: The robustness of CLIPNAL. The mismatch ratio is 40% for both datasets.
Robustness to OOD Ratio.
Fig. 7 shows the results for (a) CIFAR10 and (b) CIFAR100 with varying OOD ratios. The closer the model performance is to the black dashed line, the more informative the labeled ID data. CLIPNAL demonstrates robust performance, with minimal accuracy drop at the final round regardless of the OOD ratio. In CIFAR10, the accuracy of the classifier trained with ten rounds approximates that of a classifier trained on the entire dataset. This indicates that CLIPNAL minimizes cost loss and selects highly informative data.
@@
Robustness to Mismatch Ratio.
Fig. 8 shows the results for (a) CIFAR10 and (b) CIFAR100 with varying mismatch ratios. CLIPNAL shows significant accuracy improvement in the initial data collection iterations compared to other query strategies. In CIFAR10, its performance remains close to the classifier trained on the entire dataset regardless of the mismatch ratio. In CIFAR100, despite a larger number of classes, CLIPNAL maintains high accuracy compared to other strategies. Overall AL performance per the mismatch and OOD ratio can be found in Appendix. E.
Comparison with Standard AL Scenarios.
Fig. 9 demonstrates the robustness of CLIPNAL. The black dashed line represents the accuracy change during data collection assuming no OOD data, applying CONF to only ID data. Without OOD data, there is no cost loss. CLIPNAL, despite the presence of OOD data, shows accuracy convergence to the black dashed line in CIFAR10 and even surpasses it in CIFAR100. This indicates that CLIPNAL selects more informative data than standard AL assuming only ID data.

### v2

@@ 3.3 Visual Similarity Weighting
Although CLIPN shows high performance in detecting OOD data using textual information, it encounters challenges when the number of classes increases or when the OOD data domain is similar to that of ID data. We use visual information from ID data as additional weights to address these challenges. First, the average representation $\mu^{\text{ID}}_{j}$ is calculated by applying the image encoder $\phi$ to the labeled ID samples $\text{x}^{\text{ID}}_{j}$ for each class $j$.
$\mu^{\text{ID}}_{j}=\frac{1}{N_{j}^{\text{ID}}}\sum_{i=1}^{N_{j}^{\text{ID}}}{\phi(x_{i}^{\text{ID}})},$ (5)
where $N_{j}^{\text{ID}}$ is the number of labeled samples for class $j$.
Next, the visual similarity weight $s_{ij}$ for each class is calculated by computing the cosine similarity between the visual representation $\phi$ and the average representation $\mu^{\text{ID}}_{j}$, and converting it to probabilities using the softmax function.
$s_{ij}=\frac{\exp{(f_{i}\cdot\mu^{\text{ID}}_{j}/\tau)}}{\sum_{k=1}^{K}\exp{(f_{i}\cdot\mu^{\text{ID}}_{k}/\tau)}}.$ (6)
@@
$\mathbb{I}(x_{i})$ =$=\begin{cases}1,&p_{i}^{*,\text{OOD}}\geq\max_{j}{p_{ij}^{*,ID}},\\ 0,&\text{else},\end{cases}$
Figure 6: Changes in OOD detection accuracy of VLPure based on the temperature parameter. Both datasets are constructed with a 40% mismatch ratio and a 60% OOD ratio.
@@ 3.4 Self-Temperature Tuning
The temperature parameter significantly influences model confidence in predictions. The optimal temperature varies depending on the OOD detection data domain, as shown in Figure 6. Labeled ID and OOD data allow for exploring the optimal temperature parameter.
@@
$\text{Acc}(\tau)$ $=\frac{1}{N_{L}}\sum_{i}^{N_{L}}{\mathbb{I}(y_{i}^{\text{OOD}}=\hat{y}_{i}(\tau))},$
${\tau}^{*}$ $=\operatorname*{arg\,max}_{0<\tau}{\text{Acc}({\tau})},$
where $N_{L}$ is the number of labeled data $S=\{x_{i}\in S^{\text{ID}}\cup S^{\text{OOD}}\}$. $y_{i}^{\text{OOD}}$ indicates 1 if $x_{i}$ is an OOD sample and 0 otherwise.
According to Eq. 8, we apply self-temperature tuning to identify the optimal temperature $\tau^{*}$ by comparing the accuracy $Acc(\tau)$ against the ground truth, $y_{i}^{\text{OOD}}$.
3.5 VLPure-AL Process
Algorithm 1 summarizes the entire VLPure-AL process. AL involves data collection and model training. Initially, since no labeled data exists, data is randomly selected from the unlabeled pool for model training. Given that the unlabeled data contains OOD data, the proportion of OOD data in the chosen data reflects the overall distribution. Only ID data is used for model training, excluding OOD data.
Once initial training is complete, data is selected using the proposed VLPure-AL method. VLPure-AL assesses the purity of data from the unlabeled pool. First, it explores the optimal temperature parameter using labeled ID and OOD data, then applies CLIPN’s purity assessment process and visual weighting to calculate the purity of the unlabeled data. Based on the computed purity, OOD data is excluded, and the remaining data is assessed for informativeness using a standard AL query strategy. The selected data is annotated according to predefined costs and added to the labeled data. After data collection, the classifier is trained with the newly labeled data, and the model is reinitialized and retrained upon further data collection. Once the user-defined number of rounds is complete, the final classifier trained with the last labeled data is used as the final model.
Input: Labeled data $S=\{y_{i},x_{i}\in S^{\text{ID}}\cup S^{\text{OOD}}\}_{i=1}^{N_{\text{L}}}$, Unlabeled data $D_{\text{U}}=\{x_{i}\in\textbf{x}^{\text{ID}}\cup\textbf{x}^{\text{OOD}}\}_{i=1}^{N_{\text{U}}}$, Query $Q$, classifier $f^{\text{clf}}$, Image encoder $f^{\text{img}}$, Text encoder $f^{\text{text}}$, No text encoder $f^{\text{no text}}$, Temperature $\tau$, Annotation costs $C$, Number of classes in labeled data $K$, and Round $R$.
Output: $f^{\text{clf}}$
// Initial labeled data sampling
Randomly select samples $S=\{x_{i}\}_{i=1}^{C}$ from $D_{\text{U}}$.
Divide $S$ into $S^{\text{ID}}$ and $S^{\text{OOD}}$.
for $0\leq r\leq R$ do
       Initialize $f^{\text{clf}}$
      Train $f^{\text{clf}}$ on $S^{\text{ID}}$
      Find optimal temperature $\tau^{*}$ using Eq.8.
      $\hat{\textbf{x}}_{\text{U}}^{\text{ID}}\leftarrow$ Purity($D_{\text{U}},S^{\text{ID}},f^{\text{img}},f^{\text{text}},f^{\text{no text}},\tau^{*}$)
      $Q\leftarrow$ StandardAL($f^{\text{clf}},\hat{\textbf{x}}_{\text{U}}^{\text{ID}},C$)
      Make annotation on the selected $Q$
      $S^{\text{ID}}\leftarrow S^{\text{ID}}\cup\textbf{x}^{\text{ID}}$
      $S^{\text{OOD}}\leftarrow S^{\text{OOD}}\cup\textbf{x}^{\text{OOD}}$
      $D_{\text{U}}\leftarrow D_{\text{U}}-Q$
return $f^{\text{clf}}$
Function Purity($D_{\text{U}},S_{\text{ID}},f^{\text{img}},f^{\text{text}},f^{\text{no text}},\tau$):
       // Extract the features of unlabeled data
       Calculate $\textbf{z}^{\text{img}}$ of $D_{\text{U}}$ and $\textbf{z}^{\text{text}}$, $\textbf{z}^{\text{no text}}$ using Eq.1.
      // Calculate ID and OOD probability using ATD
       Calculate $p_{ij}^{\text{ID}}$ and $p_{ij}^{\text{OOD}}$ with $\tau$, $\textbf{z}^{\text{img}}$, $\textbf{z}^{\text{text}}$, and $\textbf{z}^{\text{no text}}$ using Eq.4
      // Adjust ID and OOD probability using visual similarity weights
       Calculate class representation $\phi_{j}$ of $S_{\text{ID}}$ using Eq.5.
      Calculate visual similarity weights $s_{ij}$ with $\tau$ and $\phi_{j}$ using Eq.6.
      Adjust $p_{ij}^{\text{ID}}$ and $p_{ij}^{\text{OOD}}$ to $p_{ij}^{*,\text{ID}}$ and $p_{ij}^{*,\text{OOD}}$ using 7.
      // Select predicted ID samples
       Select $\hat{\textbf{x}}_{\text{U}}^{\text{ID}}\leftarrow\{\mathbb{I}^{*}(x_{i})=0,x_{i}\in D_{\text{U}}\}$.
      return $\hat{\textbf{x}}_{\text{U}}^{\text{ID}}$
Algorithm 1 VLPure-AL Process
@@ 4.1 Experimental Setup
Baselines. To demonstrate the efficiency of VLPure-AL, we select query strategies from standard AL and open-set AL as comparison methods. For standard AL, the following strategies are selected: (1) uncertainty-based strategy, including CONF (Wang and Shang, 2014), LL (Yoo and Kweon, 2019), and (2) diversity-based strategy, including CORESET (Sener and Savarese, 2018). For open-set AL, the chosen strategies are (1) contrastive learning-based strategy MQNet (Park et al., 2022), (2) OOD detector-based strategy, including LfOSA (Ning et al., 2022) and EOAL (Safaei et al., 2024). To evaluate the effectiveness of query strategies, we include RANDOM, which selects data randomly without specific criteria, in the experiments.
Table 1: Benchmark description.
Datasets Image Size # Classes # Trainset # Testset CIFAR10 32 x 32 10 50,000 10,000 CIFAR100 32 x 32 100 50,000 10,000 Tiny-ImageNet 64 x 64 200 100,000 10,000 DomainNet real 224 x 224 345 120,906 52,041 clipart 224 x 224 345 33,525 14,604 infograph 224 x 224 345 36,023 15,582 painting 224 x 224 345 50,416 21,850 sketch 224 x 224 345 48,212 20,916 quickdraw 224 x 224 345 120,750 51,750
Table 2: The number of samples for open-set scenarios.
Mismatch Ratio OOD Ratio CIFAR10 & CIFAR100 Tiny-imageNet Initial $S_{L}$ $D_{U}$ Initial $S_{L}$ $D_{U}$ ID OOD ID OOD ID OOD ID OOD 20% 10% 450 50 9,550 1,061 900 100 19,100 2,122 20% 400 100 9,600 2,400 800 200 19,200 4,800 40% 300 200 9,700 6,467 600 400 19,400 12,933 60% 200 300 9,800 14,700 400 600 19,600 29,400 30% 10% 450 50 14,550 1,617 900 100 29,100 3,233 20% 400 100 14,600 3,650 800 200 29,200 7,300 40% 300 200 14,700 9,800 600 400 29,400 19,600 60% 200 300 14,800 22,200 400 600 29,600 44,400 40% 10% 450 50 19,550 2,172 900 100 39,100 4,344 20% 400 100 19,600 4,900 800 200 39,200 9,800 40% 300 200 19,700 13,133 600 400 39,400 26,267 60% 200 300 19,800 29,700 400 600 39,600 59,400
Table 3: The number of samples for DomainNet open-set scenarios.
OOD Ratio # ID Class DomainNet Initial $S_{L}$ $D_{U}$ ID OOD ID OOD real 50 50 0 15,440 394,342 clipart 50 50 0 3,944 405,838 infograph 50 50 0 4,957 404,825 painting 50 50 0 6,806 400,816 sketch 50 50 0 6,218 403,564 quickdraw 50 50 0 17,450 392,332
Open-set Data. We perform experiments on CIFAR10, CIFAR100 (Krizhevsky and Hinton, 2009), Tiny-ImageNet (Yao and Miller, 2015), and DomainNet (Peng et al., 2019), with detailed descriptions of the benchmarks provided in Table 1. To validate the effectiveness of open-set AL, we reconstruct the datasets by employing mismatch and OOD ratios for open-set data, with the detailed configurations provided in Table 2. The mismatch ratio represents the proportion of ID classes in the dataset (Ning et al., 2022; Safaei et al., 2024), while the OOD ratio denotes the proportion of OOD data in the unlabeled pool (Du et al., 2021; Park et al., 2022). Previous experiments by LfOSA and EOAL focused on open-set data constructed with the mismatch ratio but did not consider the OOD ratio. We incorporate both ratios to create more realistic scenarios when constructing open-set data. For instance, in CIFAR10 with a mismatch ratio of 20%, two classes are ID, and the remaining eight are OOD. The mismatch ratios are defined as 20%, 30%, and 40%. The OOD ratios are defined as 10%, 20%, 40%, and 60%. The composition of the open-set data used in the experiments is detailed in Table 2.
To simulate scenarios where different domains are included in the unlabeled data, we constructed an open-set configuration using the DomainNet dataset, with the detailed configurations provided in Table 3. Among the six domains in DomainNet, one domain is designated as in-domain, while the remaining five are treated as out-of-domain. Additionally, out of the 345 classes in each domain, 50 classes are selected as in-distribution, and the rest are considered out-of-distribution.
@@
We evaluate the robustness and effectiveness of the query strategies under various scenarios. The annotation cost $C$ for AL is 500 for CIFAR10 and CIFAR100, 1000 for Tiny-ImageNet, and 50 for DomainNet, where DomainNet includes one image per class. The initial data for training the classifier is randomly selected from the unlabeled pool. In subsequent iterations, data is selected based on the query strategy with a cost of $C$.
@@
Implementation Details. The image classifier is ResNet18 (He et al., 2016) in the experiments. The number of rounds $R$ for AL experiments is set to 10. Each query strategy experiment is repeated five times, with average results reported.
The experiments use ResNet18 (He et al., 2016) as the image classifier for the CIFAR10, CIFAR100, and Tiny-ImageNet datasets. The model is trained for 20,000 steps with a batch size of 64 for all datasets. The learning rate was set to 0.1, weight decay to 0.0005, and momentum to 0.9, with the model parameters updated using Stochastic Gradient Descent (SGD).
For the experiments on DomainNet, we use the ViT-B/16 (Dosovitskiy et al., 2021) pre-trained on ImageNet-21k. The pre-trained model weights are kept frozen, and only a linear classifier is trained. The model is trained for 100 epochs with a learning rate of 0.01, weight decay of 0.0005, and momentum of 0.9, with model parameters updated using Stochastic Gradient Descent (SGD).
The image encoder and text encoder for CLIPN were pre-trained on LAION-2B (Cherti et al., 2023), and the ”no” text encoder was pre-trained on CC3M (Sharma et al., 2018). The image encoder uses a ViT-B/16 with a patch size of $16\times 16$. For data informativeness, VLPure-AL uses CONF (Wang and Shang, 2014), one of the query strategies proposed in standard AL. Hyper-parameters for LL, MQNet, LfOSA, and EOAL are set as per the original studies (Yoo and Kweon, 2019; Park et al., 2022; Ning et al., 2022; Safaei et al., 2024). Experiments are conducted using Python 3.8.10 and PyTorch 1.14 on an Intel Xeon CPU E5-2698 v4 with a Tesla V100 GPU.
Figure 7: Comparison of precision changes in open-set AL based on OOD ratio. The mismatch ratio is set to 40%. Precision refers to the proportion of ID data in selected queries.
@@
Evaluation Metrics. Precision, the proportion of selected ID queries, is used to evaluate annotation cost loss in open-set AL (Ning et al., 2022). Higher precision indicates a higher proportion of ID data in the selected queries.
@@
AL results are assessed using Area Under Budget Curve (AUBC) (Zhan et al., 2021). $\text{AUBC}_{acc}$ measures the efficiency of AL by calculating the area under the curve of accuracy improvement over rounds. A larger area indicates greater accuracy improvement after data collection.
@@ 4.2 Open-set Robustness
Comparison of Annotation Cost Loss. We examine annotation cost loss through precision changes during data collection in Figure 7. Each round’s precision indicates the proportion of ID data among the selected data at the round. RANDOM, shown in black, selects data randomly, so the proportion of OOD data in the selected queries matches the OOD ratio in the unlabeled pool. A higher precision than RANDOM indicates a higher selection of ID data than the OOD ratio.
Figure 8: Comparison of accuracy changes in open-set AL based on OOD ratio. The mismatch ratio is set to 40%. The black dashed line represents the performance of a classifier trained on the entire dataset.
Standard AL strategies, except CORESET, show lower precision than RANDOM in all cases, particularly as the OOD ratio increases. This indicates significant cost loss when OOD data is in the unlabeled pool. In contrast, most open-set AL strategies show higher precision than RANDOM. However, MQNet, LfOSA, and EOAL do not show consistent precision across rounds. MQNet initially focuses on purity, selecting more ID data, but later prioritizes informativeness, causing precision to drop. LfOSA and EOAL initially had low precision due to the limited OOD data for training the detection model, but they improved as more OOD data was collected. Unlike existing strategies, VLPure-AL maintains high precision throughout data collection, regardless of the OOD ratio, indicating minimal annotation cost loss.
Figure 9: Comparison of accuracy changes in open-set AL based on mismatch ratio. The OOD ratio is set to 60%. The black dashed line represents the performance of a classifier trained on the entire dataset.
Robustness to OOD Ratio. Figure 8 shows the results for (a) CIFAR10, (b) CIFAR100, and (c) Tiny-ImageNet with varying OOD ratios. The closer the model performance is to the black dashed line, the more informative the labeled ID data. VLPure-AL demonstrates robust performance, with minimal accuracy drop at the final round regardless of the OOD ratio. In CIFAR10, the accuracy of the classifier trained with ten rounds approximates that of a classifier trained on the entire dataset. This indicates that VLPure-AL minimizes cost loss and selects highly informative data.
@@
Robustness to Mismatch Ratio. Figure 9 shows the results for (a) CIFAR10, (b) CIFAR100, and (c) Tiny-ImageNet with varying mismatch ratios. VLPure-AL shows significant accuracy improvement in the initial data collection iterations compared to other query strategies. CIFAR10’s performance remains close to the classifier trained on the entire dataset regardless of the mismatch ratio. In CIFAR100 and Tiny-ImageNet, despite a larger number of classes, VLPure-AL maintains high accuracy compared to other strategies.
Table 4: $\text{AUBC}_{acc}$(%) performance of query strategies based on the mismatch and OOD ratio.
[…]
Open-set Robustness. Table 4 summarizes the $\text{AUBC}_{acc}$ results for all open-set scenarios. MQNet shows high $\text{AUBC}_{acc}$ in CIFAR10 with high OOD and mismatch ratios but does not outperform standard AL in all cases. In CIFAR100 and Tiny-ImageNet, RANDOM often outperforms other strategies. However, VLPure-AL consistently shows higher $\text{AUBC}_{acc}$ than existing methods, indicating significant accuracy improvement from the initial round.
Table 5: $\text{AUBC}_{acc}$(%) and Last ACC(%) performance of query strategies on DomainNet. ”w/ domain” indicates that domain information is provided in the prompt format ”A photo of {domain} {class}.”
[…]
Domain Robustness. Table 5 shows the experimental results for Open-set AL when different classes and domains are mixed as OOD data. In the case of Open-set DomainNet, the unlabeled pool contains more complex OOD samples compared to CIFAR10, CIFAR100, and Tiny-ImageNet. As a result, Open-set AL methods generally outperform not only RANDOM but also Standard AL. For VLPure-AL, a significant performance improvement is observed when text information providing domain context (e.g., ”A photo of {domain} {class}”) is used compared to cases where such information is not provided. Moreover, except for the quickdraw domain, VLPure-AL consistently achieves higher $\text{AUBC}_{acc}$ and Last ACC in almost all cases, highlighting the advantage of incorporating domain information depending on the scenario.
Figure 10: The robustness of VLPure-AL. The mismatch ratio is 40% for both datasets.
Comparison with Standard AL Scenarios. Figure 10 demonstrates the robustness of VLPure-AL. The black dashed line represents the accuracy change during data collection assuming no OOD data, applying CONF to only ID data. Without OOD data, there is no cost loss. VLPure-AL, despite the presence of OOD data, shows accuracy convergence to the black dashed line in CIFAR10 and even surpasses it in CIFAR100. This indicates that VLPure-AL selects more informative data than standard AL assuming only ID data.
Table 6: Comparison of zero-shot, full-shot, and VLPure-AL performance. Zero-shot represents the performance of CLIP, while full-shot refers to the performance achieved by training only on ID samples. Both full-shot and the classifier used for active learning utilize CLIP’s visual encoder.
Datasets Mismatch Ratio Accuracy (%) zero-shot round 0 round 1 round 2 round 3 full-shot CIFAR10 20% 99.76 99.76 99.75 99.80 99.83 99.77 30% 99.06 99.15 99.21 99.43 99.37 99.32 40% 98.54 98.75 98.88 99.14 99.16 99.04 CIFAR100 20% 90.90 91.19 93.04 93.80 93.97 94.29 30% 87.67 85.96 87.59 89.53 90.35 91.41 40% 84.31 80.62 83.14 85.33 85.88 89.31 Tiny-ImageNet 20% 82.79 81.62 84.85 86.65 86.99 88.70 30% 79.89 76.30 80.17 82.05 83.14 85.90 40% 77.98 71.92 74.99 77.44 78.50 83.45
Comparison AL with zero-shot and full-shot. Table 6 shows the difference between zero-shot and full-shot performance when applying active learning to open-set data. Full-shot refers to the performance achieved by training solely on ID samples. The classifier used for active learning employed the visual encoder from CLIP, the same one used for the zero-shot results, to allow for a direct comparison. While zero-shot performance using CLIP is generally strong compared to full-shot, the results demonstrate that collecting data through active learning can surpass zero-shot performance, especially when the mismatch ratio is small, requiring only a small number of samples. Furthermore, VLPure-AL rapidly reaches full-shot performance within just four rounds, even in open-set data scenarios.
4.3 Ablation Studies